OpenStudy (anonymous):

Simplify by removing factors of 1 n^2-4/(n+2)^2

OpenStudy (amistre64):

(n+2)(n-2) --------- = (n-2)/(n+2) (n+2)(n+2)

OpenStudy (anonymous):

\[{n^2-4 \over (n+2)^2}={(n-2)(n+2) \over (n+2)(n+2)}={n-2 \over n+2}\]

OpenStudy (anonymous):

so that is the simplified form?

OpenStudy (anonymous):

Yep.